If you invest $100 in the U.K market for one year with a forward cover when the interest rate and exchange rates are as given below, the annualized rate of return from the investment is:

            Spot rate = $1.3050/£; forward rate = $1.3000/£, U.K. interest rate = 4%
Answer #1

Converted $100 into pounds = 100/1.3050

= Pounds 76.6283525

Invest for one year and get 76.6283525*1.04

= Pounds 79.69348659

Convert back into USD = 79.69348659*1.3000

= $103.6015

Hence, return from investment = (103.6015-100)/100

= 3.6015%